An aluminum sphere has a diameter of 15.6 cm. What is its mass? (Refer to the table in the densities of common substances for aluminum)

1
Expert's answer
2021-11-18T12:21:22-0500

Let's first find the volume of the aluminum sphere:


"V=\dfrac{4}{3}\pi r^3=\dfrac{4}{3}\pi\times(0.078\ m)^3=0.002\ m^3."

Finally, we can find the mass of the aluminum sphere:


"m=\rho V=2700\ \dfrac{kg}{m^3}\times0.002\ m^3=5.4\ kg."